OpenStudy (anonymous):

Evaluate: 9 − 3 · 2 + 4

OpenStudy (anonymous):

9 - 6 + 4 = 7

OpenStudy (anonymous):

PEMDAS or BODMAS order of operations, is how polycosm got their answer

OpenStudy (anonymous):

ok PEMDAS it is the order you calculate things P- Parenthesis ( do you the things inside the parenthesis first while still following order of operations) E- Exponents - you calculate the exponents (square roots are considered exponents and everything within a square root is considered within a parenthesis) MD- multiplication/ division - you do these operations next ( you will learn later than division is the same as multiplying 1 over the number or multiplying by a fraction AS- addition/ subtraction- these are done last so lets look at your question 9 − 3 · 2 + 4 no parenthesis or exponents, so we go to multiplication/ division we multiply 3*2 to get 6 9-6+4 no more multiplication/division we now add/ subtract 9-6=3 so we get 3+4 and 3+4 =7 so the answer is 7